# DyCodeEval: Dynamic Benchmarking of Reasoning Capabilities in Code Large Language Models Under Data Contamination

### v1

@@ 4.4 Programming Problem Diversity
To evaluate the diversity of the generated programming problems, we conduct two experiments: one for external diversity and one for internal diversity. External diversity quantifies the dissimilarity between the generated and seed problems, while internal diversity measures the diversity within each problem-generation method across trials. We use two metrics: BLEU-4 to measure syntactical diversity and cosine similarity of the prompt’s semantic embedding to measure semantic diversity. For semantic embedding, we use the GPT-2 model to obtain the embedding of each natural language prompt. Moreover, we also consider PPM (Chen et al., 2024) and a series of robustness-based mutation (Wang et al., 2023), such as token replacement, insert blank lines, as our comparison baseline.
@@ 4.6 Impact of Foundation LLM
5 Application
Leveraging the dynamic nature of our method, we propose a new metric, DivPass, to address the limitations of the current gold standard, Pass@K. Unlike Pass@K, which generates candidate solutions for a fixed problem and evaluates correctness, our approach creates semantic variants of a seed problem. These variants preserve the original problem’s complexity by modifying only the description while maintaining the same underlying algorithmic abstraction. Additionally, the variants expand the input space beyond that of Pass@K, making it more challenging to achieve full coverage. As a result, DivPass provides a more rigorous assessment of code LLMs’ reasoning abilities, particularly under potential data contamination. Compared to Pass@K, which evaluates solutions within a fixed problem context, DivPass introduces contextual variations during benchmarking. This allows it to better distinguish whether a model is merely memorizing problem context or genuinely reasoning to solve it.
@@ 5 Application
To demonstrate the advantages of DivPass, we compare it against Pass@K on both contaminated and wild models, with for evaluation. The results are presented in Table 5 and Table 5. From the results in Table 5, we observe that when the model is trained on leaked data, the static metric Pass@K fails to accurately reflect the model’s reasoning capabilities, with all Pass@K scores rising to very high levels (e.g., from 0.82 to 0.89). In contrast, our dynamic metric DivPass @K shows a slight decrease rather than a significant increase, highlighting the sensitivity of DivPass to data contamination. When comparing Pass@K and DivPass @K on models that were not specifically trained on the leaked dataset, both metrics show consistency in benchmarking code LLMs. Based on these observations, we conclude that our dynamic metric, DivPass, effectively reflects the reasoning capabilities of code LLMs, even under data contamination. Moreover, DivPass @K aligns with static benchmarking metrics when there is no data contamination.
@@ 6 Conclusion
Assessing the overall capabilities of LLMs is crucial in order to maintain the reliability and safety of model usage in society. Data contamination, however, raises an issue by causing inflated accuracy in model evaluation. Our work proposes a new benchmark DyCodeEval, designed to accurately measure a model’s true capabilities allowing us to deepen our understanding of them.
@@ Appendix A Proof of Theorem
The total number of possible distinct outcomes is , the size of the random space, let Since each of the samples must **not** match , and they are drawn independently, the exact probability is:
@@ Appendix C Prompt Templates & Scenario Examples
Scenario Examples
Real World Domain | Scenario |
Banking | Predictive and Personal Detection in Financial Transactions |
AI | Automated Customer Service Chatbots |
Transportation | Predictive Maintenance for Fleet Vehicles |
Banking | Personalized Financial Risk Assessment |
Healthcare | Early Disease Prediction and Prevention |
Education | Adaptive Learning Path Recommendation |
Social Networking | Content Relevance and Personalization |
Education | Adaptive Learning Path Personalization |
Social Networking | Content Recommendation and Engagement Optimization |
Scenario Proposer
Suggest a real-world scenarios that could provide meaningful context. The scenarios should be general yet practical, covering areas such as {s1}, {s2}, {s3}, {s4}, {s5}, and other practical areas. Return each scenario on a separate line for clarity and just return the scenario without another reasoning steps.
Scenario example:
example
{exp}
example
Suggested Real-world Scenarios:
scenario
Content Generator
I have a natural language problem description, a description of the input types, and a real-world scenario. For each variable in the problem, suggest a meaningful context based on the provided scenario.
# Problem Description:
problem_description
{pb}
problem_description
# Input Types:
input_types
{var}
input_types
# Real-world Scenario:
scenario
{scenario}
scenario
# Context for each Variable:
context
Prompt Rewriter
Given an seed programming problem description, a selected real-world scenario, and contextualized input variables, rewrite the original problem to fit the new context. The rewritten problem description should preserve the original problem’s complexity and constraints while making it relevant to the given scenario. Ensure the new prompt is clear, concise, and maintains solution applicability. Just return the new rewritten problem description without any other texts.
# Problem Description:
{pb}
# Real World Scenario:
{scenario}
# Contextualized Input Variables:
{input_variables}
# Rewritten Problem Description
Validator 1
Assess whether the two given natural language instructions have the same meaning. Provide your answer as either ’Yes’ or ’No’ only. # Instruction A: {inst_a} # Instruction B: {inst_b} # Your Answer:
Validator 2
Does the following code solve the problem described in the Instruction. Provide your answer as either ’Yes’ or ’No’ only. # Instruction: {inst_a} # Code Solution: {code} # Your Answer:
@@ Appendix D Human Verification
In order to add an extra level of validation between the original and generated prompts, we employ manual human verification. Given a benchmarked dataset and DyCodeEval-generated questions for each problem, we randomly pick N pairs of questions (each pair consists of a benchmarked problem and a newly generated question). Each pair of questions is then examined by two graduated student level student to determine if the core algorithm and complexity of the question is preserved, if there are inconsistency between these students, they discuss the case until the agreement is reached. Based on the radomly sampled N pairs of questions, the consistent rate is around 95%.

### v2

@@ 4.6 Impact of Foundation LLM
5 Dynamic Evaluation Metrics
Leveraging the dynamic nature of our method, we propose a new metric, DyPass, to address the limitations of the current gold standard, Pass@K. Unlike Pass@K, which generates candidate solutions for a fixed problem prompt and evaluates the correctness, our approach creates semantic prompt variants of a seed problem. These prompt variants preserve the complexity of the original problem by modifying only the description while maintaining the same underlying algorithmic abstraction. Furthermore, prompt variants expand the input space beyond that of Pass@K, making it more challenging to achieve full coverage. As a result, DyPass provides a more rigorous assessment of code LLMs’ reasoning abilities, particularly under potential data contamination. Compared to Pass@K, which evaluates solutions within a fixed problem context, DyPass introduces contextual variations during benchmarking. This allows it to better distinguish whether a model is merely memorizing the problem context or genuinely reasoning to solve it.
@@ 5 Dynamic Evaluation Metrics
To demonstrate the advantages of DyPass, we compare it against Pass@K on both contaminated and in-the-wild models, with for evaluation. The results are presented in Table 5 and Table 5. From the results in Table 5, we observe that when the model is trained on leaked data, the static metric Pass@K fails to accurately reflect the model’s reasoning capabilities, with all Pass@K scores rising to very high levels (e.g., from 0.82 to 0.89). In contrast, our dynamic metric DyPass @K shows a slight decrease rather than a significant increase, highlighting the sensitivity of DyPass to data contamination. When comparing Pass@K and DyPass @K on models that were not specifically trained on the leaked dataset, both metrics show consistency in benchmarking code LLMs. Based on these observations, we conclude that our dynamic metric, DyPass, effectively reflects the reasoning capabilities of code LLMs, even under data contamination. Moreover, DyPass @K aligns with static benchmarking metrics when there is no data contamination.
@@ 6 Conclusion
Acknowledgements
This work was supported in part by CCF 2313055, CCF 2107405, CAREER 2025082, and FAI: 2040961. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors.
@@ Acknowledgements
Assessing the overall capabilities of large language models (LLMs) is essential for ensuring their reliable and safe deployment in society. However, data contamination can inflate evaluation accuracy, obscuring a model’s true performance. To address this, we propose a new benchmarking method, DyCodeEval, which enables more accurate measurement of LLM capabilities and provides deeper insights into their behavior.
@@ Appendix A Proof of Theorem
The total number of possible distinct outcomes is , the size of the random space, let Since each of the samples must not match , and they are drawn independently, the exact probability is:
@@ Appendix C Prompt Templates & Scenario Examples
C.1 Template for Scenario Proposer Agent
C.2 Example for Scenario Proposer Agent
C.3 Prompt for Context Generator Agent
C.4 Example for Context Generator Agent
C.5 Prompt for Prompt Rewriter Agent
C.6 Example for Prompt Rewriter Agent
C.7 Prompt for Validation Agent 1
C.8 Example for Validation Agent 1
C.9 Prompt for Validation Agent 2
C.10 Example for Validation Agent 2
@@ Appendix D Human Verification
To add an additional layer of validation between the original and DyCodeEval-generated prompts, we perform a small-scale manual verification. Given a benchmark dataset and the corresponding generated questions, we randomly sample problem pairs from each dataset (60 in total), where each pair consists of a benchmark problem and its generated variant. Each pair is independently reviewed by two graduate-level students to assess whether the core algorithm and complexity are preserved. In cases of disagreement, the reviewers discuss the discrepancies until consensus is reached. Out of the 60 reviewed pairs, the annotators initially disagreed on three but were able to resolve all disagreements through discussion, resulting in an overall agreement rate of 95%.